OpenStudy (anonymous):

find the perimeter of a triangle with vertices (2,4) and (-1,3) and (2,-3)

OpenStudy (anonymous):

You'll need the formulat for the distance between two points.\[\sqrt{(x _{1}-x _{2})^{2}+(y _{1}-y _{2})^{2}}\] Now you plug in the values from two of your points. For example (2,4) and (-1,3). 2-(-1) and 4-3 will give you. 3 and 1. \[\sqrt{3^{2}+1^{2}}\] That means you will have 9 + 1 = 10 \[\sqrt{10}\] that is the length of one side of your triangle. Do the same thing for the other sides of your triangle and add them together.